# How Numerical Precision Affects Mathematical Reasoning Capabilities of LLMs

### v1

@@ 14 Covert $p^{m}$-adic $\tilde{\bm{o}}$ to $p$-adic ${\bm{o}}$:
Block 4. This block of the Transformer executes the fourth line of Algorithm 2, i.e., decompose $c_{i,j}t_{i,j}$ as $b_{i,j}p^{m}+q_{i,j}$. Notice that $b_{i,j}\leq i$, thus $b_{i,j}$ will not exceed the absolute positional of the current token. We can set the query, key and value as follows:
  • Query: ${\bm{q}}_{i,j}=\left(-(c_{i,j}t_{i,j}+\frac{1}{2})^{2},2p^{m}(c_{i,j}t_{i,j}+\frac{1}{2}),-p^{2m}\right)$.
  • Key: ${\bm{k}}_{i^{\prime},j^{\prime}}=\left(1,\text{ape}_{i^{\prime},j^{\prime}}-\frac{1}{2},(\text{ape}_{i^{\prime},j^{\prime}}-\frac{1}{2})^{2}\right)$.
  • Value: ${\bm{v}}_{i^{\prime},j^{\prime}}=\text{ape}_{i^{\prime},j^{\prime}}$.
These values can be calculated using Lemma C.1. Thus
$\langle{\bm{q}}_{i,j},{\bm{k}}_{i^{\prime},j^{\prime}}\rangle=-\left[c_{i,j}t_{i,j}-\left(\text{ape}_{i^{\prime},j^{\prime}}-\frac{1}{2}\right)p^{m}+\frac{1}{2}\right]^{2},$
@@
which implies that $|\langle{\bm{q}}_{i,j},{\bm{k}}_{i^{\prime},j^{\prime}}\rangle|\leq\left(\frac{p^{m}-1}{2}\right)^{2}$ if $\text{ape}_{i^{\prime},j^{\prime}}=\lfloor\frac{c_{i,j}t_{i,j}}{p^{m}}\rfloor$, and $\langle{\bm{q}}_{i,j},{\bm{k}}_{i^{\prime},j^{\prime}}\rangle\leq-\left(\frac{p^{m}+1}{2}\right)^{2}$ otherwise. Using Lemma C.7,
$c=\frac{(p^{m}+1)^{2}-(p^{m}-1)^{2}}{(p^{m}-1)^{2}}\geq\frac{4}{p^{m}},$
@@
thus $1/c=O(p^{m})=O(k)$. This implies we can set the query, key and value, such that the attention output is $\lfloor\frac{c_{i,j}t_{i,j}}{p^{m}}\rfloor=b_{i,j}$. Finally, we calculate $q_{i,j}$ by $q_{i,j}=c_{i,j}t_{i,j}-p^{m}b_{i,j}$ using the subsequent MLP. The new embeddings generated in this block is ${\bm{e}}^{4}=(b,q)$.
@@
Block 5. This block of the Transformer computes $q_{w+1}+b_{w}$ for $s_{w}$. Recall that, $s_{w}=c_{i,mw}t_{i,mw}$ where $i$ is the maximize value such that the length of ${\bm{a}}_{i}$ is greater than $mw$. We want to calculate those values at the corresponding positions.
@@
First, we use attention layer to COPY $q_{w+1}$ for token ${\bm{a}}_{i}$ defined above. Notice that we can always get the correct value since the position with correct value of $s_{w+1}$ is always in front of that with correct value of $s_{k}$. To finish this, we use the attention mechanism to COPY from the position of the value $s_{w+1}$. This can be implemented by setting query, key, value and $r$ in Section C.2:
@@
We can calculate the values required in query or key by previous MLPs using Lemma C.1. Thus $\langle{\bm{q}}_{i,j},{\bm{k}}_{i^{\prime},j^{\prime}}\rangle=-\left(\lfloor\frac{j^{\prime}}{m}\rfloor-\lfloor\frac{j}{m}\rfloor-1\right)^{2}-(j^{\prime}\operatorname{mod}m)^{2}$, which implies that $\langle{\bm{q}}_{i,j},{\bm{k}}_{i^{\prime},j^{\prime}}\rangle=0$ if $\lfloor\frac{j^{\prime}}{m}\rfloor=\lfloor\frac{j}{m}\rfloor+1$ and $j^{\prime}\operatorname{mod}m=0$, and $\langle{\bm{q}}_{i,j},{\bm{k}}_{i^{\prime},j^{\prime}}\rangle\leq-1$ otherwise.
[…]
${\bm{1}}_{q_{w+1}+b_{w}\geq p^{m}}=\text{ReLU}[q_{w+1}+b_{w}-(p^{m}-1)]-\text{ReLU}[q_{w+1}+b_{w}-p^{m}]$
and Lemma C.2.
To sum up, the embeddings generated in this block are:
  • Positions with the correct value of $s_{w}$:
[…]
  • Other positions: ${\bm{e}}^{6}=(-1,-1,-1,-1,-1,-1,-1)$. (This can be implemented by filtering out the unfeasible values using Lemma C.3).
Block 6. This block of the Transformer first calculates the following values for the positions with the correct value of $s_{w}$:
  • The minimum $w_{1}\geq w$ such that ${\bm{1}}_{q_{w+1}+b_{w}\geq p^{m}}=1$.
  • The minimum $w_{2}\geq w$ such that ${\bm{1}}_{q_{w+1}+b_{w}\leq p^{m}-2}=1$.
Both are standard COPY operation, thus can be implemented by Lemma C.5. To make sure the value of $w_{1},w_{2}$ is valid (i.e., the existence of $w_{1},w_{2}$), we can COPY ${\bm{1}}_{q_{w_{1}+1}+b_{w_{1}}\geq p^{m}},{\bm{1}}_{q_{w_{2}+1}+b_{w_{2}}\leq p^{m}-2}$ and check whether they are $1$, and filter out the invalid values by MLP using Lemma C.3. The embeddings generated in this block are:
  • Positions with the correct value of $s_{w}$: ${\bm{e}}^{6}=(w_{1},w_{2})$.
  • Other positions: ${\bm{e}}^{6}=(-1,-1)$. (This can be implemented by filtering out the unfeasible values using Lemma C.3).
@@
For any integers $n$, $l$, and $k$, there exists a polynomial $f(n)$ such that a logarithmic-precision Transformer with constant depth and hidden dimensions $f(n)$ can generate the correct output for any input on the $\operatorname{MUL}(n,l,p)$ task.
Here, we first describe an algorithm to perform $\operatorname{Mul}(n,l,p)$ (Algorithm 3) and prove the correctness of Algorithm 3. Then, we construct a Transformer with the configurations in Theorem 5.3 capable for simulating Algorithm 3.
@@ 1 $m=\lceil\log_{p}n\rceil+1$;
2Compute the product of each pair of bits: $d_{i,j}=a_{i}b_{j}$;
@@ 9       $g_{i}={\bm{1}}_{q_{i}+b_{i-1}\geq p^{m}-2}$;
13 Compute the $p^{m}$-adic outcome $\tilde{\bm{o}}$: $\tilde{o}_{i}=(q_{i}+b_{i-1}+c_{i-1})\operatorname{mod}p^{m}$ for $i=0,\cdots,\lfloor(2n-1)/m\rfloor$;
@@ 14 Covert $p^{m}$-adic $\tilde{\bm{o}}$ to $p$-adic ${\bm{o}}$:
Algorithm 3 $\operatorname{Mul}(n,l,p)$ Algorithm
Lemma E.2 (An algorithm to perform $\operatorname{Mul}(n,l,p)$).
@@ F.2 Model Training
We tried 3 different seeds and select the maximum accuracy for each task. All the experiments were conducted on a single NVIDIA GeForce RTX 4090 GPU over a duration of two weeks.
@@ F.3 Integer Addition Results
F.4 Prompt For LLM
Detailed prompts for the three elementary arithmetic tasks are listed in the Tables 4 and 5 and generation configuration can be found in the Table 3.
@@ F.4 Prompt For LLM
Please directly calculate the following arithmetic expression in base base with the following format:
@@
Please directly calculate the following arithmetic expression in base base.

### v2

@@ 14 Covert $p^{m}$-adic $\tilde{\bm{o}}$ to $p$-adic ${\bm{o}}$:
Block 4. This block of the Transformer corresponds to the fourth step in Algorithm 2, which decomposes $c_{i,j}t_{i,j}$ as $b_{i,j}p^{m}+q_{i,j}$. It is important to observe that $b_{i,j}\leq i$, meaning that $b_{i,j}$ does not exceed the absolute positional index of the current token. To achieve this decomposition, we define the query, key, and value as follows:
  • Query: ${\bm{q}}_{i,j}=\left(-(c_{i,j}t_{i,j}+\frac{1}{2})^{2},2p^{m}(c_{i,j}t_{i,j}+\frac{1}{2}),-p^{2m}\right)$
  • Key: ${\bm{k}}_{i^{\prime},j^{\prime}}=\left(1,\text{ape}_{i^{\prime},j^{\prime}}-\frac{1}{2},(\text{ape}_{i^{\prime},j^{\prime}}-\frac{1}{2})^{2}\right)$
  • Value: ${\bm{v}}_{i^{\prime},j^{\prime}}=\text{ape}_{i^{\prime},j^{\prime}}$
The above components can be computed using Lemma C.1. Consequently, the inner product of the query and key is given by:
$\langle{\bm{q}}_{i,j},{\bm{k}}_{i^{\prime},j^{\prime}}\rangle=-\left[c_{i,j}t_{i,j}-\left(\text{ape}_{i^{\prime},j^{\prime}}-\frac{1}{2}\right)p^{m}+\frac{1}{2}\right]^{2}.$
@@
This expression implies that $|\langle{\bm{q}}_{i,j},{\bm{k}}_{i^{\prime},j^{\prime}}\rangle|\leq\left(\frac{p^{m}-1}{2}\right)^{2}$ if $\text{ape}_{i^{\prime},j^{\prime}}=\lfloor\frac{c_{i,j}t_{i,j}}{p^{m}}\rfloor$, and $\langle{\bm{q}}_{i,j},{\bm{k}}_{i^{\prime},j^{\prime}}\rangle\leq-\left(\frac{p^{m}+1}{2}\right)^{2}$ otherwise. By applying Lemma C.7, we obtain:
$c=\frac{(p^{m}+1)^{2}-(p^{m}-1)^{2}}{(p^{m}-1)^{2}}\geq\frac{4}{p^{m}}.$
@@
From this, it follows that $1/c=O(p^{m})=O(k)$. Hence, we can design the query, key, and value such that the attention output satisfies $\lfloor\frac{c_{i,j}t_{i,j}}{p^{m}}\rfloor=b_{i,j}$. Finally, $q_{i,j}$ can be computed as $q_{i,j}=c_{i,j}t_{i,j}-p^{m}b_{i,j}$ using the subsequent MLP. The embeddings generated by this block are thus given by ${\bm{e}}^{4}=(b,q)$.
@@
Block 5. This block of the Transformer computes $q_{w+1}+b_{w}$ for $s_{w}$. Recall that $s_{w}=c_{i,mw}t_{i,mw}$, where $i$ is the largest index such that the length of ${\bm{a}}_{i}$ is greater than $mw$. The goal is to compute these values at their corresponding positions.
@@
First, we use the attention mechanism to copy $q_{w+1}$ for the token ${\bm{a}}_{i}$ defined above. Note that it is always possible to retrieve the correct value because the position associated with the correct value of $s_{w+1}$ precedes that of $s_{k}$. To achieve this, we utilize the attention mechanism to copy from the position containing the value $s_{w+1}$. This is implemented by appropriately configuring the query, key, value, and $r$ as described in Section C.2:
@@
The values required for the query or key can be computed using previous MLPs, as shown in Lemma C.1. The dot product $\langle{\bm{q}}_{i,j},{\bm{k}}_{i^{\prime},j^{\prime}}\rangle$ evaluates to
$\langle{\bm{q}}_{i,j},{\bm{k}}_{i^{\prime},j^{\prime}}\rangle=-\left(\lfloor\frac{j^{\prime}}{m}\rfloor-\lfloor\frac{j}{m}\rfloor-1\right)^{2}-(j^{\prime}\operatorname{mod}m)^{2}.$
This implies $\langle{\bm{q}}_{i,j},{\bm{k}}_{i^{\prime},j^{\prime}}\rangle=0$ if $\lfloor\frac{j^{\prime}}{m}\rfloor=\lfloor\frac{j}{m}\rfloor+1$ and $j^{\prime}\operatorname{mod}m=0$, and $\langle{\bm{q}}_{i,j},{\bm{k}}_{i^{\prime},j^{\prime}}\rangle\leq-1$ otherwise.
Using Lemma C.5, one attention head suffices to copy the values $q_{i^{\prime},j^{\prime}}$ and $\lfloor\frac{j^{\prime}}{m}\rfloor$ from the last token satisfying the conditions. Consequently, the first dimension of the attention output equals $q_{w+1}$ if an input number with length greater than $m(w+1)$ exists, as required. Otherwise, $q_{w+1}$ should be zero, and the attention output remains undefined. These two cases can be distinguished by inspecting the second dimension of the attention output: if no input number has a length greater than $m(w+1)$, the second dimension of the attention output is at most $\lfloor\frac{j}{m}\rfloor$. Using Lemma C.3, we can identify these cases and set $q_{w+1}=0$ when necessary.
Finally, a subsequent MLP computes the correct value of $q_{w+1}+b_{w}$ for $s_{w}$. Additionally, this MLP calculates the indicators ${\bm{1}}_{q_{w+1}+b_{w}\geq p^{m}}$, ${\bm{1}}_{q_{w+1}+b_{w}\leq p^{m}-2}$, ${\bm{1}}_{b_{w}\geq p^{m}}$, and ${\bm{1}}_{b_{w}\leq p^{m}-2}$ using the formulation:
${\bm{1}}_{q_{w+1}+b_{w}\geq p^{m}}=\text{ReLU}[q_{w+1}+b_{w}-(p^{m}-1)]-\text{ReLU}[q_{w+1}+b_{w}-p^{m}],$
as described in Lemma C.2.
To summarize, the embeddings generated in this block are as follows:
  • For positions with the correct value of $s_{w}$:
    ${\bm{e}}^{5}=(q_{w+1}+b_{w},b_{w},{\bm{1}}_{q_{w+1}+b_{w}\geq p^{m}},{\bm{1}}_{q_{w+1}+b_{w}\leq p^{m}-2},{\bm{1}}_{b_{w}\geq p^{m}},{\bm{1}}_{b_{w}\leq p^{m}-2},w).$
  • For all other positions:
    ${\bm{e}}^{5}=(-1,-1,-1,-1,-1,-1,-1).$
    This can be achieved by filtering out infeasible values using Lemma C.3.
Block 6. This block of the Transformer computes the following values for positions with the correct value of $s_{w}$:
  • The smallest $w_{1}\geq w$ such that ${\bm{1}}_{q_{w+1}+b_{w}\geq p^{m}}=1$.
  • The smallest $w_{2}\geq w$ such that ${\bm{1}}_{q_{w+1}+b_{w}\leq p^{m}-2}=1$.
Both calculations rely on the standard COPY operation, which can be implemented using Lemma C.5. To ensure the validity of $w_{1}$ and $w_{2}$ (i.e., the existence of such indices), we COPY the values ${\bm{1}}_{q_{w_{1}+1}+b_{w_{1}}\geq p^{m}}$ and ${\bm{1}}_{q_{w_{2}+1}+b_{w_{2}}\leq p^{m}-2}$, verifying that they equal $1$. Invalid values can then be filtered out using an MLP, as described in Lemma C.3.
The embeddings generated in this block are as follows:
  • For positions with the correct value of $s_{w}$: ${\bm{e}}^{6}=(w_{1},w_{2})$.
  • For all other positions: ${\bm{e}}^{6}=(-1,-1)$. (This can be achieved by filtering infeasible values using Lemma C.3.)
@@
Fix integers $p\geq 2$ and $c\in\mathbb{N}^{*}$. Consider the tokenizer ${\bm{T}}_{c}$ defined in Eq. 1 for processing the input and output sequences. For any integers $n$ and $l\leq 2n$, there exists a logarithmic-precision Transformer with constant depth (independent of $n$ and $k$) and hidden dimensions $O(n^{2})$ that can generate the correct output for any input on the $\operatorname{MUL}_{p}(n,l)$ task.
Here, we first describe an algorithm to perform $\operatorname{MUL}_{p}(n,l)$ (Algorithm 3) and prove the correctness of Algorithm 3. Then, we construct a Transformer with the configurations in Theorem 5.3 capable for simulating Algorithm 3.
@@ 1 $m=\lceil\log_{p}n\rceil+1$;
2 Compute the product of each pair of bits: $d_{i,j}=a_{i}b_{j}$;
@@ 9       $g_{i}={\bm{1}}_{q_{i}+b_{i-1}\geq p^{m}-2}$;
13
14Compute the $p^{m}$-adic outcome $\tilde{\bm{o}}$: $\tilde{o}_{i}=(q_{i}+b_{i-1}+c_{i-1})\operatorname{mod}p^{m}$ for $i=0,\cdots,\lfloor(2n-1)/m\rfloor$;
@@ 15 Covert $p^{m}$-adic $\tilde{\bm{o}}$ to $p$-adic ${\bm{o}}$:
Algorithm 3 $\operatorname{MUL}_{p}(n,l)$ Algorithm
Lemma E.2 (An algorithm to perform $\operatorname{MUL}_{p}(n,l)$).
@@ F.2 Model Training
The experiments were conducted on a single NVIDIA GeForce RTX 4090 GPU over a duration of two weeks, investigating the differences in performance between standard precision and low precision operations. To avoid some unexpected issues of hardware, we also conduct the same experiments on NVIDIA A100 GPUs, and the results are consistent with the results on NVIDIA GeForce RTX 4090 GPU. We try 3 different seeds and select the maximum accuracy for each task.
The model configuration in our experiments is presented in Table 2, and the training configuration is presented in Table 3.
Model Configuration
Model Depth $\{3,5\}$
Hidden Dimension 256
Attention Heads 4
Positional Embeddings RoPE
Activation NewGeLU
Table 2: Model Configuration for Transformer in Experiments.
Training Configuration
Epochs 1
Learning Rate 1e-3
Optimizer AdamW
$\beta_{1}$ 0.9
$\beta_{2}$ 0.999
Weight Decay 0.01
Learning Rate Scheduler Cosine Scheduler with Warmup
[…]
Table 3: Training Configuration in Experiments.
@@ F.3 Integer Addition Results
F.4 Fine-tuing Configuration, Generation Configuration, and Prompt For LLM
The fine-tuning configuration and generation configuration for LLMs is listed in Tables 6 and 5. The detailed prompts for the three elementary arithmetic tasks are listed in the Tables 7 and 8 and generation configuration can be found in the Table 5.
@@ F.4 Fine-tuing Configuration, Generation Configuration, and Prompt For LLM
Fine-tuning Configuration
Rank 8
Scaling Factor 16
Dropout Rate 0.05
Epochs 1
Learning Rate 2e-4
Optimizer AdamW
$\beta_{1}$ 0.9
$\beta_{2}$ 0.999
Weight Decay 0.01
Learning Rate Scheduler Cosine Scheduler with Warmup
Warmup Ratio 0.1
[…]
@@
Please directly calculate the following arithmetic expression in base <base> with the following format:
@@
Please directly calculate the following arithmetic expression in base <base>.
@@
F.5 Reference Results for LLMs
We also provide the results of GPT-4o and GPT-4o-mini as a baseline for these arithmetic tasks base-10 for reference. The results are presented in Table 9.
Task Length GPT-4o GPT-4o-mini
Addition of 2 numbers 1 100.0% 100.0%
4 99.9% 98.8%
7 97.5% 51.4%
10 96.3% 46.0%
13 93.3% 44.0%
Addition of 3 numbers 1 100.0% 100.0%
3 99.8% 99.6%
5 98.9% 73.4%
7 69.2% 9.1%
9 38.5% 5.8%
Addition of 5 numbers 1 100.0% 100.0%
2 100.0% 99.4%
3 100.0% 89.5%
4 88.4% 31.1%
5 86.8% 24.7%
Multiplication of 2 numbers 1 100.0% 100.0%
2 100.0% 97.5%
3 76.6% 44.7%
4 21.5% 7.6%
5 4.1% 0.7%
Table 9: The Performance of GPT-4o and GPT-4o-mini on the arithmetic tasks.